Volcán Barva (01.01.2013 to 03.31.2013)

You are here

PDF versionPDF version
Year: 
2013
Quarter: 
January - March
Communications activities: 
  • The Site Manager gave a talk to students of OTS tropical biology course on January 2013 about TEAM network.
  • News from our animal pictures photo-trapped in this 2013 sampling period in La Selva was published in the OTS webpage and the e-canopy newsletter.
Unusual Events at the TEAM Site: 

A new record of jaguar and Tapir within La Selva biological station have been an interesting sighting. The first record of jaguar was caught on 2005.

New Species: 

A new tree species for La Selva was recorded in our annual census of vegetation within the 1 ha plot located at 100 masl. This species (Pradosia atroviolacea) belongs to the family Sapotacea.

Protocol Activities: 
  • The climate data from January 1st up to March 4th were collected and submitted in the new data entry form version 3.0.
  • The annual census of trees, palms and lianas ≥10 cm diameter in the 1 ha plot located at 100, 200, 400, 900, 2800, masl was completed up to date. More than 3000 stems have been measured.
  • We have completed the terrestrial vertebrates monitoring for the sampling period in array 1 and 2. (low and middle elevation). 40 camera traps were deployed. Interesting sightings of 16 mammals were recorded. 4000 pictures were processed on the deskteam.
  • All data and cards from Array 1 were already sent to CI office in Washington.
Protocol Problems: 
  • On January we replaced all batteries for a big one of 150Ah. After we install this battery, we have not had problems downloading the data from the datalogger. The datalogger and all sensors are functioning well. We prefer use a small computer directly for download the data instead of cards that sometimes malfunctioning. Using the computer we assure the transfer of the data. In addition, we took advantage of this field trip to remove one of the RH/T sensors that was malfunctioning due the head of the sensor was rusted.
  • I have uploaded all information required for the new sensors and the old sensors removed for calibration, but the new sensor does not appear on the climate maintenance webpage. I mean, when I am doing the climate maintenance for these Rh/T sensor, the serial number does not appear.
  • After end the terrestrial sampling period for array 1 and 2, we noted 6 cameras malfunctioning. Basically, the cameras are misfiring due probably a problem with the LED sensor. We need to replace these cameras for the next year sampling period.
Schedule Problems: 

There were not shedule problems in this sampling period.

Logistical Problems: 

We currently have some mechanic problems with the motorcycle and we could not collect the climate data on March.

Suggestions: 
  • Regarding climate data uploaded, I have noted some errors in the climate status webpage. In some dates the system put missing data and it is pending an explanation. However climate data for that dates have already uploaded.
  • I would like to know step by step how to export the terrestrial vertebrates data from the deskteam.